How many millimeters of pure acid must be added to 120 milliliters of a 30% acid solution to obtain a 50% solution?
Answer by checkley77(12569)

.30*120+x=.50(120+x)
36+x=60+.5x
x-.5x=60-36
.5x=24
x=24/.5
x=48 ml of pure acid is used.
Proof:
.30*120+48=.5(120+48)
36+48=.5*168
84=84